1. A 2D inviscid flow field is represented by the velocity potential function: ° = Ax + Bx2 – By2. Where A = 1m/s, B = 15-7,help
Answer #1

GIVEN:

\tiny \phi = Ax+Bx2-By2

A= 1 m/s, B = 1 S-1

\tiny \rho = 1.2 kg/m3

TO FIND :

a) VELOCITY FIELD

b) FLOW IS IRROTATIONAL

c) FLOW IS INCOMPRESSIBLE

d) EXPRESSION OF STREAM FUNCTION

e)PRESSURE DIFFERENCE BETWEEN THE POINTS(0,0) AND (1,2)

SOLUTION:

a) VELOCITY POTENTIAL

Function of space such that negative derivative of this function with respect to a given direction will give the component of velocity in that direction.

T= ; 9 V= ay ;  W = z

therefore

T=

(Ax + Bx2 - By2) U= ax

u = -(A+2Bx) .............(1)

substituting A and B IN .....(1)

U = (1+2(1)X)

U = -(1+2X)

subs points(0,0) and (1,2)

U(0,0) = -(1+ 2(0))

U(0,0) = -1 m/s

similarly,

U(1,2) = - (1+ 2(1))

U(1,2) = -(1+2 )= -3 m/s

U(1,2)= -3 m/s

and

9 V= ay

V= O(Ax + Bx2 - By2) ay

v = -(-2y)

V= 2By

V =2y

V(0,0) = 2(0)

V(0,0) = 0 m/s

V(1,2) = 2(2)

V(1,2) = 4 m/s

the velocity equation is V = ui + vj + wk

since it is 2D (w=0)

V(0,0) = -i at origin is along negative x axis

at (0,0) the velocity is V1 =\small \left ( \sqrt{-1} \right )^{2} = 1 m/s

V(1,2) = -3i +4j   

at (1,2) the velocity is V2 = \small \sqrt{\left ( -3^{2} + 4^{2}\right )} = (9+16)

V2 =\small \sqrt{\left ( 25}\right )}  

V2 = 5 m/s

b) FLOW IS IRROTATIONAL

condition for flow is  irrotational (ند 0)

= 1 (ди 2\дх ди ду,

u = -(A+2Bx) ; v = 2y

av ax 0 ;ди ду о

(0-0) II

ند 0

therefore, the flow is irrotational

c) FLOW IS INCOMPRESSIBLE

condition for incompressible flow  2) 2) + = 0 дX2 ду?.........................is that continuty equation should be satisfied

=   22(Ax + Bx? – By?) ax2 + (Ax + Bx? – By?) ay2

= д(A + 2Bх) дх + д-2By) ду

= 2B-2B

= 0

since continuity equation is satisfied , it is incompressible flow.

d) EXPRESSION OF STREAM FUNCTION

for 2D flow

=V ax ;   \frac{\partial \psi }{\partial y} = -u ;  

д дх 2y..................................................................(eqn A)

др || 2удх

V = 2yx + c..............................................................(eqn B)

similarly

\frac{\partial \psi }{\partial y} = -u

(A + 2Bx) ду...............................................(eqn c)

the constant of integration c is a function of y in .....eqn B

V = 2yx + c differentiating with respect to y

д ay 2x + дс ay...........................................................(eqn E)

subsituting eqn C in eqn E

(A+2Bx) - 2x= \frac{\partial c}{\partial y}

substituting A and B values

1 + 2x – 2x = да ду

де 1 ду

1ду - Јә =

= с

subs c= y in eqn B

U = 2yx + y

therefore the stream function is U = 2yx + y

e)PRESSURE DIFFERENCE BETWEEN THE POINTS(0,0) AND (1,2)

\Delta p = \frac{1}{2}\rho \left ( V_{2}^{2}-V_{1}^{2} \right )

\Delta p = \frac{1}{2}*1.2* \left (5^{2}-1^{2} \right )

\Delta p = \frac{1}{2}*1.2* \left (24 \right )

\Delta p = 14.4 N/m^{2}

RESULT :

  

a) VELOCITY FIELD is V(0,0) = -i at origin is along negative x axis and V(1,2) = -3i +4j   

b) FLOW IS IRROTATIONAL , since   ند 0

c) FLOW IS INCOMPRESSIBLE , since it satisfies bernoulli equation  2) 2) + = 0 дX2 ду?

d) EXPRESSION OF STREAM FUNCTION is U = 2yx + y

e)PRESSURE DIFFERENCE BETWEEN THE POINTS(0,0) AND (1,2) is \Delta p = 14.4 N/m^{2} .